One mole of an ideal gas (group 18) at 27 ℃ and 1 atm expands adiabatically from a volume of 10 dm3 to 20 dm3. Calculate:
(i) q
(ii) w
(iii) ΔU
(iv) ΔH
(v) Final temperature of the gas.
1
Expert's answer
2020-12-06T13:52:23-0500

T= 27°C= 27+273

P= 1atm

V1= 10dm3

V2= 20dm3


1. q= 0, for an adiabatic expansion, heat is neither gained nor lost by the system.


2. w= -P(V2-V1)

w= -1(20-10)

w= -10L.atm

1L.atm= 101.3J

-10L.atm= -10 x 101.3J= -1013J


3. ∆U = q+w

Since q= 0 (adiabatic)

∆U=w

∆U=-1013J


4. ∆H=∆U+P∆V

∆H=-1013+1(20-10)

∆H= -1013+10

∆H= -1003J


5. V1= 10dm3 T1= 300K

V2= 20dm3 T2= ?

V1/T1 = V2/T2

T2= V2T1/V1 = 20x300/10

= 600K = 327°C